Jason, Eric, and Kylie have a combined age of 38. Jason is 4 years older than kylie and eric is 2 years younger than kylie what is the age of each

Found 2 solutions by ewatrrr, TimothyLamb:
Answer by ewatrrr(24785)

x Kylie's age
Jason + Eric +Kylie = 38
(x+4) + (x-2) + x = 38
3x = 36
x = 12, Kylie's age. Jason 16 and Eric is 10
And...checking
sum = 38

Answer by TimothyLamb(4379)

x = jason age now
y = eric age now
z = kylie age now
---
x + y + z = 38
x = z + 4
y = z - 2
---
put the system of linear equations into standard form
---
x + y + z = 38
x - z = 4
y - z = -2
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = jason age now = 16
y = eric age now = 10
z = kylie age now = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---